OpenStudy (anonymous):

8x squared - 11x=-3 I need to find the number of real solutions.

OpenStudy (anonymous):

like this? 8x^2 - 11x = -31 if so, we will need to use the quadratic equation: \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

do you know how to pick out the values for a, b, and c from the equation you posted?

OpenStudy (anonymous):

its not 31 but its equals 3

OpenStudy (anonymous):

8x^2 - 11x = -3 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8x^2 - 11x = -3 becoms 8x^2 - 11x + 3= 0 can you pick out the a, b, and c values of your equation looking at this similar form? ax^2 + bx + c = 0 a=? b=? c=?